The bulk modulus of material is 3.72 × 1011 N/m2. What fractional change in volume does a piece of this material undergo when it is subjected to a bulk stress increase of 8.46x107 N/m2? Assume that the force is applied uniformly over the surface


1
Expert's answer
2021-03-31T09:40:38-0400
"\frac{\Delta V}{V}=\frac{\Delta p}{B}\\\frac{\Delta V}{V}=\frac{8.46\cdot10^{7}}{3.72\cdot10^{11}}=2.27\cdot10^{-4}"